# Revisiting Reweighted Risk for Calibration: AURC, Focal Loss, and Inverse Focal Loss

### v2

@@ Abstract
Several variants of reweighted risk functionals, such as focal losss, inverse focal loss, and the Area Under the Risk–Coverage Curve (AURC), have been proposed in the literature and claims have been made in relation to their calibration properties. However, focal loss and inverse focal loss propose vastly different weighting schemes. In this paper, we revisit a broad class of weighted risk functions commonly used in deep learning and establish a principled connection between these reweighting schemes and calibration errors. We show that minimizing calibration error is closely linked to the selective classification paradigm and demonstrate that optimizing a regularized variant of the AURC naturally leads to improved calibration. This regularized AURC shares a similar reweighting strategy with inverse focal loss, lending support to the idea that focal loss is less principled when calibration is a desired outcome. Direct AURC optimization offers greater flexibility through the choice of confidence score functions (CSFs). To enable gradient-based optimization, we introduce a differentiable formulation of the regularized AURC using the SoftRank technique. Empirical evaluations demonstrate that our AURC-based loss achieves competitive class-wise calibration performance across a range of datasets and model architectures.
@@ 1 Introduction
Over the past decade, advances in deep learning have enabled neural networks to achieve, and in some cases surpass, human-level performance on tasks ranging from image recognition [19] to natural language understanding [1, 33]. However, superior accuracy alone does not guarantee that the predicted confidence scores faithfully reflect true outcome probabilities – a phenomenon known as miscalibration [17]. Miscalibration can lead to critical errors in high-stakes applications where trustworthy uncertainty estimates are essential. To address and quantify this gap, the research community has introduced a range of calibration error metrics [5, 17, 30, 41] to evaluate the discrepancy between predicted confidences and observed outcome frequencies. Recent work has explored both trainable [24, 31, 26, 4, 41] and post-hoc [17, 10, 29, 44, 36] approaches for neural network calibration, achieving impressive reductions in terms of those calibration errors. Among trainable methods, focal loss [32]—which emphasizes hard examples during training—has gained popularity for its calibration benefits. Conversely, inverse focal loss [43], which focuses on the easy examples more, has also been shown to improve calibration in certain settings. This raises a key question: Which weighting scheme is most effective for improving model calibration?
We aim to answer this question by examining the relationship between selective classification [13] and calibration error. We provide a theoretical foundation that links calibration error minimization to the principles of selective classification. Building on this connection, we propose a loss function derived from the Area Under the Risk-Coverage Curve (AURC) [14], originally introduced in the context of selective classification, and demonstrate that it adopts a reweighting scheme similar to that of the inverse focal loss. To enable gradient-based optimization, we leverage the SoftRank [3] method to approximate AURC in a differentiable manner. This adaptation ensures that our AURC-based loss remains differentiable with respect to any choice of confidence score function (CSF).
@@
1 We establish a theoretical connection between calibration error minimization and selective classification by analyzing the structure of calibration error and its relation to the “abstain” strategy at the core of selective classification.
2 Motivated by this connection, we propose a differentiable loss function based on the AURC metric. Although it adopts a reweighting strategy similar to that of inverse focal loss, it offers greater flexibility in the choice of CSFs.
3 We demonstrate through extensive experiments that our proposed AURC-based loss achieves competitive class-wise calibration performance across multiple datasets and network architectures, outperforming existing calibration-aware loss functions.
@@ 2 Related work
Post-hoc calibration methods address miscalibration by learning a mapping function that transforms model outputs into calibrated posterior probabilities. Approaches like Platt scaling [40], temperature scaling (TS) [17], and local temperature scaling [10] typically preserve classification accuracy by introducing a scalar temperature parameter. In contrast, other probabilistic methods—including Beta calibration [28], Dirichlet calibration [29], Gaussian processes [44], and the Concrete distribution [11]—may achieve better calibration but often at the cost of reduced classification accuracy. Notably, Dirichlet calibration generalizes Beta calibration to multi-class settings. Ma and Blaschko [36] and Liu et al. [35] proposed post-hoc calibration methods that adjust predicted probabilities based on correctness, using binary classifiers to either distinguish between correct and incorrect predictions or to directly modulate confidence accordingly. However, these approaches often involve a trade-off between calibration and predictive accuracy. We aim to develop a trainable calibration method that improves calibration performance without significantly compromising predictive accuracy.
@@
Trainable calibration methods seek to enhance the calibration of deep neural networks by explicitly modifying the training objective. One class of approaches introduces differentiable surrogate losses that directly approximate the expected calibration error, as demonstrated in [24, 31, 26, 4, 41]. Popordanoska et al. [41] proposed $\mathrm{ECE}^{\mathrm{KDE}}$, a differentiable estimator of calibration error, and introduced the KDE-XE objective, which augments cross-entropy (XE) with an ECE-KDE regularizer to balance accuracy and calibration. Alternatively, several studies replace the standard cross-entropy loss with calibration-aware objectives, including mean squared error loss [23], inverse focal loss [43], and focal loss [32]. Focal loss, in particular, has been empirically shown to improve calibration by mitigating overconfident predictions [37, 6, 25]. Nevertheless, Wang et al. [43] argue that the regularization effect of focal loss may suppress important information about sample difficulty, thereby limiting the effectiveness of subsequent post-hoc calibration techniques. To address this trade-off, several variants of focal loss have been proposed, such as Adaptive Focal Loss [15] and Dual Focal Loss [42]. Distinct from these prior approaches, we propose a loss function based on the selective classification framework. While selective classification is not originally intended for calibration, we show that its underlying principle is fundamentally aligned with minimizing calibration error.
@@
Selective classification augments a standard classifier with a reject option, enabling an explicit trade-off between selective risk and coverage. This technique mitigates selective risk by withholding low-confidence predictions [13, 14, 9, 12]. The Area Under the Risk-Coverage Curve (AURC) and its normalized variant, Excess-AURC (E-AURC) [14], are among the most widely used evaluation metrics for selective classification systems. These metrics compute the risk associated with accepted predictions across different confidence thresholds. Zhou et al. [47] showed that AURC can be interpreted as a weighted risk function, where the weights are determined by the ranking induced by a CSF $g$. This work is motivated by the goal of establishing a theoretical and empirical connection between selective classification and model calibration. We empirically show that minimizing a regularized AURC objective yields strong calibration performance across a range of architectures and datasets. While the underlying weighting scheme is similar to that of the inverse focal loss, our approach offers substantially greater flexibility in the choice of CSFs, making it more adaptable to different modeling scenarios.
@@ 3.1 Weighted Risk Functions
Focal loss reduces to cross-entropy when $\gamma=0$. More generally, it can be interpreted as a reweighted version of cross-entropy, where each instance is assigned an importance weight $\boldsymbol{w}(\boldsymbol{p})=(1-\boldsymbol{p})^{\gamma}\in\mathbb{R}^{k}$. The weight depends on both the predictive probability $p_{y^{\prime}}$ and the focusing parameter $\gamma$: samples predicted with high confidence are downweighted, while harder or less confident samples are emphasized when $\gamma\geq 1$.
@@
where $p_{y^{\prime}}$ denotes the predicted probability for the true class label $y^{\prime}$.
@@
(b) Loss with Normalized Weights
Figure 1: Comparison of weight and loss behavior with respect to prediction probability $p$. We normalize the weights in both focal loss and inverse focal loss for fair comparison.
In contrast to focal loss, which downweights confident predictions, inverse focal loss assigns greater weight to high-confidence samples. We analytically derive the gradient of inverse focal loss (App. 7.3) and show empirically that certain values of $\gamma$ may lead to undesirable behavior during training (see Fig. 5 in appendix). Surprisingly, despite their fundamentally opposite reweighting schemes, both strategies have been reported to improve classifier calibration.
@@ 3.2 Selective classification
Selective classification pairs a classifier $f$ with a confidence scoring function $g:\mathcal{X}\rightarrow[0,1]$, which estimates the reliability of its predictions. A selective classifier, defined as the pair $(f,g)$, makes a prediction only when the confidence score exceeds a threshold $\tau\in\mathbb{R}$, and abstains otherwise:
$(f,g)(\boldsymbol{x}):=\begin{cases}f(\boldsymbol{x})&\text{if }g(\boldsymbol{x})\geq\tau,\\ \text{"abstain"}&\text{otherwise}.\end{cases}$ (4)
@@
$R(f,g):=\frac{\mathbb{E}\left[\ell(f(x),y)\cdot\mathbb{I}\left(g(x)\geq\tau\right)\right]}{\mathbb{E}\left[\mathbb{I}\left(g(x)\geq\tau\right)\right]}.$ (5)
where $\mathbb{I}$ is the indicator function and the denominator represents the Coverage. The threshold $\tau$ determines the trade-off between selective risk and coverage. Rather than evaluating performance at a single threshold, the AURC [14] assesses classifier robustness across all possible thresholds, as induced by the distribution of confidence scores. This metric can equivalently be expressed as a weighted risk function [47], as described below.
@@
Let $\boldsymbol{p}=f(\boldsymbol{x})$ be a predicted probability vector produced by the model $f$ with fixed input $\boldsymbol{x}$. Denote with $\boldsymbol{p}^{\prime}=f\left(\boldsymbol{x}^{\prime}\right)$ the output random variable with input random variable $\boldsymbol{x}^{\prime}$. The cumulative distribution function (CDF) of the CSF $g(\boldsymbol{p})$ is defined as:
[…]
@@
$\text{AURC}_{g}(f)=-\mathbb{E}\left[\ln\left(1-G(\boldsymbol{p})\right)\cdot\ell(\boldsymbol{p},\boldsymbol{y})\right].$ (7)
This formulation reveals that AURC can be interpreted as a reweighted risk, where each sample is weighted by $w(\boldsymbol{p})=-\ln(1-G(\boldsymbol{p}))$, a function that depends on the CSF. Common choices for the CSFs include the Maximum Softmax Probability (MSP), i.e., $g(\boldsymbol{p})=\max_{i=1}^{k}p_{i}$, or the negative entropy of the logits prior to the softmax layer as described in Table 3.
@@ 3.3 Calibration Errors
A model $f:\mathcal{X}\rightarrow\Delta^{k}$ is calibrated if and only if $f(\boldsymbol{x})=\mathbb{E}[\boldsymbol{y}\mid f(\boldsymbol{x})]$ almost surely based on the input random variable $\mathbf{x}$.
@@
$\text{CE}_{\rho}(f)=\left(\mathbb{E}\left[\|f(\boldsymbol{x})-\mathbb{E}[\boldsymbol{y}\mid f(\boldsymbol{x})]\|_{\rho}^{\rho}\right]\right)^{1/\rho},$ (8)
where the expectation is taken with respect to $\boldsymbol{x}$.
When $\rho=1$ and for binarized outcomes, this quantity is referred to as the Expected Calibration Error (ECE) [38, 17], i.e.,
$\operatorname{ECE}\left(f\right)=\mathbb{E}\left[\lvert\max_{c}f_{c}(\boldsymbol{x})-\mathbb{P}(\boldsymbol{y}=\arg\max_{c}f_{c}(\boldsymbol{x})\mid\max_{c}f_{c}(\boldsymbol{x}))\rvert\right].$ (9)
In practice, the true value of $\text{CE}_{\rho}(\cdot)$ is unobservable, as it depends on the unknown conditional distribution $\mathbb{P}(\boldsymbol{y}\mid f\left(\boldsymbol{x}\right))$. To address this, Naeini et al. [38] proposed an empirical estimator based on a finite dataset. This estimator employs a binning scheme $B$ to approximate the true difference between the predicted confidence and the conditional probability $\mathbb{E}[\boldsymbol{y}\mid f(\boldsymbol{x})]$ by computing empirical averages within each bin. We denote this estimator as $\widehat{\text{ECE}}_{B}$.
@@
$\widehat{\text{ECE}}_{B}(f)=\sum_{b=1}^{m}\frac{n_{b}}{n}\left|\text{acc}_{b}(f)-\text{conf}_{b}(f)\right|$ (10)
@@
To avoid potential confusion, we emphasize here the empirical estimator of ECE measures the top-label calibration error instead of the calibration error w.r.t. a fixed class. The class-wise calibration error (cwECE) [30, 29] measures the expected calibration errors across all classes.
@@
$\text{cwECE}_{\rho}(f)=\Big{(}\frac{1}{k}\sum_{c=1}^{k}\mathbb{E}\left[\left|f_{c}(\boldsymbol{x})-\mathbb{P}(y^{\prime}=c\mid f_{c}(\boldsymbol{x}))\right|^{\rho}\right]\Big{)}^{1/\rho},$ (11)
@@
We follow related works [39, 18] in defining this metric as the average calibration error across classes. Similar to ECE, it can be estimated empirically using a binning scheme $B$, resulting in a bin-based estimator $\widehat{\text{cwECE}}_{B}(f)$.
Definition 3.8 (Binned estimator $\widehat{\text{cwECE}}_{B}$).
$\widehat{\mathrm{cwECE}}_{B}(f)$ $=\frac{1}{k}\sum_{c=1}^{k}\frac{1}{n}\sum_{b=1}^{m_{c}}n_{c,b}\Bigl{|}\underbrace{\frac{1}{n_{c,b}}\!\!\sum_{i\in I_{c,b}}\!\!\mathbb{I}(y_{i}^{\prime}=c)}_{\text{acc}_{c,b}}-\underbrace{\frac{1}{n_{c,b}}\!\!\sum_{i\in I_{c,b}}\!f_{c}(\boldsymbol{x}_{i})}_{\text{conf}_{c,b}}\Bigr{|}$ (12)
where $n_{c,b}$ is the number of samples in bin $I_{c,b}$, and $n$ is the total number of samples.

### v3

# Revisiting Reweighted Risk for Calibration: AURC, Focal and Inverse Focal Loss

@@ Abstract
Several variants of reweighted risk functionals, such as focal loss, inverse focal loss, and the Area Under the Risk–Coverage Curve (AURC), have been proposed for improving model calibration, yet their theoretical connections to calibration errors remain unclear. In this paper, we revisit a broad class of weighted risk functions commonly used in deep learning and establish a principled connection between calibration error and selective classification. We show that minimizing calibration error is closely linked to the selective classification paradigm and demonstrate that optimizing selective risk in low-confidence region naturally leads to improved calibration. This loss shares a similar reweighting strategy with dual focal loss but offers greater flexibility through the choice of confidence score functions (CSFs). Our approach uses a bin-based cumulative distribution function (CDF) approximation, enabling efficient gradient-based optimization without requiring expensive sorting and achieving $\mathcal{O}(nK)$ complexity. Empirical evaluations demonstrate that our method achieves competitive calibration performance across a range of datasets and model architectures.
@@ 1 Introduction
Over the past decade, advances in deep learning have enabled neural networks to achieve, and in some cases surpass, human-level performance on tasks ranging from image recognition (Haggenmüller et al.,, 2021) to natural language understanding (Achiam et al.,, 2023; Liu et al., 2024a, ). However, achieving high accuracy does not necessarily ensure that the predicted confidence scores align with the true probabilities of outcomes, a phenomenon known as miscalibration (Guo et al.,, 2017). Miscalibration can lead to critical errors in high‐stakes applications where trustworthy uncertainty estimates are essential. To address and quantify this gap, the research community has introduced a range of calibration error metrics (Brier,, 1950; Guo et al.,, 2017; Kumar et al.,, 2019; Popordanoska et al.,, 2022) to evaluate the discrepancy between predicted confidences and observed outcome frequencies. Recent work has explored both trainable (Karandikar et al.,, 2021; Kumar et al.,, 2018; Krishnan and Tickoo,, 2020; Bohdal et al.,, 2023; Popordanoska et al.,, 2022) and post-hoc (Guo et al.,, 2017; Ding et al.,, 2021; Kull et al.,, 2019; Wenger et al.,, 2020; Ma and Blaschko,, 2021) approaches for neural network calibration, achieving impressive reductions in terms of those calibration errors. Among trainable methods, focal loss (Lin et al.,, 2017)—which emphasizes hard examples during training—has gained popularity for its calibration benefits. Conversely, inverse focal loss (Wang et al.,, 2021), which focuses on the easy examples more, has also been shown to improve calibration in certain settings. This raises a key question: Which weighting scheme is most effective for improving model calibration?
We address this question by examining the relationship between selective classification (Geifman and El-Yaniv,, 2017) and calibration error. We provide a theoretical foundation that links calibration error minimization to the principles of selective classification, showing that optimal calibration strategies naturally emerge from selective weighting based on confidence rankings. Building on this connection, we propose a loss function similar to the AURC (Geifman et al.,, 2019), but focused specifically on selective risk in low-confidence regions that require careful calibration. Unlike AURC, our method employs a bin-based CDF approximation that enables efficient gradient-based optimization with $\mathcal{O}(nK)$ complexity. This adaptation ensures computational scalability while maintaining differentiability with respect to any choice of CSFs.
@@
  • We establish a theoretical connection between calibration error minimization and selective classification by analyzing the structure of calibration error and its relation to the “abstain” strategy at the core of selective classification.
  • Motivated by this connection, we propose a computationally efficient, differentiable loss function based on selective classification principles, using a bin-based CDF approximation that avoids expensive sorting operations while maintaining gradient flow for any CSF choice.
  • Through extensive experiments, we show that our proposed method delivers competitive calibration performance across various datasets and network architectures, surpassing the majority of existing calibration-aware loss functions.
@@ 2 Related work
Post-hoc calibration methods address miscalibration by learning a mapping function that transforms model outputs into calibrated posterior probabilities. Approaches like Platt scaling (Platt et al.,, 1999), temperature scaling (TS) (Guo et al.,, 2017), and local temperature scaling (Ding et al.,, 2021) typically preserve classification accuracy by introducing a scalar temperature parameter. In contrast, other probabilistic methods—including Beta calibration (Kull et al.,, 2017), Dirichlet calibration (Kull et al.,, 2019), Gaussian processes (Wenger et al.,, 2020), and the Concrete distribution (Esaki et al.,, 2024)—may achieve better calibration but often at the cost of reduced classification accuracy. Notably, Dirichlet calibration generalizes Beta calibration to multi-class settings. Ma and Blaschko, (2021) and Liu et al., 2024b proposed post-hoc calibration methods that adjust predicted probabilities based on correctness, using binary classifiers to either distinguish between correct and incorrect predictions or to directly modulate confidence accordingly. However, these approaches often involve a trade-off between calibration and predictive accuracy. We aim to develop a trainable calibration method that improves calibration performance without significantly compromising predictive accuracy.
@@
Trainable calibration methods seek to enhance the calibration of deep neural networks by explicitly modifying the training objective. One class of approaches introduces differentiable surrogate losses that directly approximate the expected calibration error, as demonstrated in (Karandikar et al.,, 2021; Kumar et al.,, 2018; Krishnan and Tickoo,, 2020; Bohdal et al.,, 2023; Popordanoska et al.,, 2022). Popordanoska et al., (2022) proposed $\mathrm{ECE}^{\mathrm{KDE}}$, a differentiable estimator of calibration error, and introduced the KDE-XE objective, which augments cross-entropy (XE) with an ECE-KDE regularizer to balance accuracy and calibration. Alternatively, several studies replace the standard cross-entropy loss with calibration-aware objectives, including mean squared error loss (Hui,, 2020), inverse focal loss (Wang et al.,, 2021), and focal loss (Lin et al.,, 2017). Focal loss, in particular, has been empirically shown to improve calibration by mitigating overconfident predictions (Mukhoti et al.,, 2020; Charoenphakdee et al.,, 2021; Komisarenko and Kull,, 2024). Nevertheless, Wang et al., (2021) argue that the regularization effect of focal loss might suppress crucial information about sample difficulty, which in turn limits the effectiveness of subsequent post-hoc calibration techniques. To address this trade-off, several variants of focal loss have been proposed, such as Adaptive Focal Loss (Ghosh et al.,, 2022) and Dual Focal Loss (Tao et al.,, 2023). Distinct from these prior approaches, we propose a loss function based on the selective classification framework. While selective classification is not originally intended for calibration, we show that its underlying principle is fundamentally aligned with minimizing calibration error.
@@
Selective classification augments a standard classifier with a reject option, enabling an explicit trade-off between selective risk and coverage. This technique mitigates selective risk by withholding low-confidence predictions (Geifman and El-Yaniv,, 2017; Geifman et al.,, 2019; Ding et al.,, 2020; Galil et al.,, 2023). The Area Under the Risk-Coverage Curve (AURC) and its normalized variant, Excess-AURC (E-AURC) (Geifman et al.,, 2019), are among the most widely used evaluation metrics for selective classification systems. These metrics compute the risk associated with accepted predictions across different confidence thresholds. Zhou et al., (2025) showed that AURC can be interpreted as a weighted risk function, where the weights are determined by the ranking induced by a CSF. Our work is motivated by the goal of establishing a theoretical and empirical connection between selective classification and model calibration. We propose a differentiable loss function that emphasize the selective risks in low-confidence region where calibration is most critical. Unlike AURC computation that requires expensive sorting operations, our approach uses a bin-based CDF approximation to achieve computational efficiency while maintaining differentiability, making it practical for large-scale deep learning applications.
@@ 3.1 Weighted Risk Functions
Focal loss reduces to cross-entropy when $\gamma=0$. More generally, it can be interpreted as a reweighted version of cross-entropy, where each instance is assigned an importance weight $\boldsymbol{w}(\boldsymbol{p})=(1-\boldsymbol{p})^{\gamma}\in\mathbb{R}^{k}$. The weight depends on both the predictive probability $p_{y^{\prime}}$ and the focusing parameter $\gamma$: samples predicted with high confidence are downweighted, while harder or low confident samples are emphasized when $\gamma\geq 1$.
@@
where $p_{y^{\prime}}$ is the predicted probability for true class $y^{\prime}$.
@@
(b) Loss
Figure 1: Comparison of weight and loss (with normalized weights) behavior with respect to prediction probability pp. We normalize the weights in both focal loss and inverse focal loss for fair comparison.
In contrast to focal loss, which downweights confident predictions, inverse focal loss assigns greater weight to high-confidence samples. We analyze the gradient of inverse focal loss and show empirically that certain values of $\gamma$ may lead to undesirable behavior during training (see Fig. 4 in Appendix C). Surprisingly, despite their fundamentally opposite reweighting schemes, both strategies have been reported to improve classifier calibration.
@@ 3.2 Selective classification
Selective classification pairs a classifier $f$ with a confidence scoring function $g:\mathcal{X}\rightarrow\mathbb{R}$, which estimates the reliability of its predictions. A selective classifier, defined as the pair $(f,g)$, makes a prediction only when the confidence score exceeds a threshold $\tau\in\mathbb{R}$, and abstains otherwise:
$(f,g,\tau)(\boldsymbol{x}):=\begin{cases}f(\boldsymbol{x})&\text{if }g(\boldsymbol{x})\geq\tau,\\ \text{"abstain"}&\text{otherwise}.\end{cases}$ (4)
@@
$\mathcal{R}(f,g,\tau):=\frac{\mathbb{E}\left[\ell(f(\boldsymbol{x}),\boldsymbol{y})\cdot\mathbb{I}\left(g(\boldsymbol{x})\geq\tau\right)\right]}{\mathbb{E}\left[\mathbb{I}\left(g(\boldsymbol{x})\geq\tau\right)\right]}.$ (5)
where $\mathbb{I}$ is the indicator function and the denominator represents the Coverage. The threshold $\tau$ determines the trade-off between selective risk and coverage.
Rather than evaluating performance at a single threshold, the AURC (Geifman et al.,, 2019) assesses classifier robustness across all possible thresholds, as induced by the distribution of confidence scores. This metric can equivalently be expressed as a weighted risk function (Zhou et al.,, 2025), as described below.
@@
Let $f(\boldsymbol{x})$ be a predicted probability vector for fixed input $\boldsymbol{x}$. Denote $\boldsymbol{x}^{\prime}$ as random input variable, the cumulative distribution function (CDF) of the CSF $g(\boldsymbol{x})$ is defined as:
$G(\boldsymbol{x})=\Pr\left(g(\boldsymbol{x}^{\prime})\leq g(\boldsymbol{x})\right)=\int\mathbb{I}\left[g(\boldsymbol{x}^{\prime})\leq g(\boldsymbol{x})\right]\,d\mathbb{P}(\boldsymbol{x}^{\prime}).$ (6)
@@
$\text{AURC}(f)=-\mathbb{E}\left[\ln\left(1-G(\boldsymbol{x})\right)\cdot\ell(f(\boldsymbol{x}),\boldsymbol{y})\right].$
This formulation shows that AURC can be interpreted as a reweighted risk, utilizing a weighting scheme similar to that of the inverse focal loss (see Fig. 1). In this scheme, each sample is weighted by a function that depends on the CSF. Common choices for the CSFs include Maximum Softmax Probability (MSP) (Hendrycks and Gimpel,, 2022), Negative Entropy (Liu et al.,, 2020), and Softmax Margin (Belghazi and Lopez-Paz,, 2021) as described in Appendix Table 1.
@@ 3.3 Calibration Errors
A model $f:\mathcal{X}\rightarrow\Delta^{k}$ is calibrated if and only if $f(\boldsymbol{x})=\mathbb{E}[\boldsymbol{y}\mid f(\boldsymbol{x})]$ almost surely based on the input random variable $\boldsymbol{x}$.
@@
$\text{CE}_{\rho}(f)=\left(\mathbb{E}\left[\|f(\boldsymbol{x})-\mathbb{E}[\boldsymbol{y}\mid f(\boldsymbol{x})]\|_{\rho}^{\rho}\right]\right)^{1/\rho},$
where the expectation is taken with respect to $\mathbb{P}(\boldsymbol{x},\boldsymbol{y})$.
When $\rho=1$ and for binarized outcomes, this quantity is referred to as the Expected Calibration Error (ECE) (Naeini et al.,, 2015; Guo et al.,, 2017). In practice, the true value of $\text{CE}_{\rho}(\cdot)$ is unobservable, as it depends on the unknown conditional distribution $\mathbb{P}(\boldsymbol{y}\mid f\left(\boldsymbol{x}\right))$. To address this, Naeini et al., (2015) proposed an empirical estimator that uses a binning scheme $B$ to approximate the calibration error, denoted by $\widehat{\text{ECE}}_{B}$.
@@
$\widehat{\text{ECE}}_{B}(f)=\sum_{b=1}^{m}\frac{n_{b}}{n}\left|\text{acc}_{b}(f)-\text{conf}_{b}(f)\right|$
@@
To avoid potential confusion, we emphasize here the empirical estimator of ECE measures the top-label calibration error instead of the calibration error w.r.t. a fixed class. The class-wise calibration error (cwECE) (Kumar et al.,, 2019; Kull et al.,, 2019) measures the expected calibration errors across all classes.
@@
$\text{cwECE}_{\rho}(f)=\Big(\frac{1}{k}\sum_{c=1}^{k}\mathbb{E}\left[\left|f_{c}(\boldsymbol{x})-\mathbb{P}(y^{\prime}=c\mid f_{c}(\boldsymbol{x}))\right|^{\rho}\right]\Big)^{1/\rho}$
@@
We follow related works (Panchenko et al.,, 2022; Gupta and Ramdas,, 2022) in defining this metric as the average calibration error across classes. Similar to ECE, it can be estimated empirically using a binning scheme $B$, resulting in a bin-based estimator $\widehat{\text{cwECE}}_{B}(f)$ (see Appendix A.1).